8:00: start of activity, telescope at horizon, WG and DP go up to install the camera to measure the LAS pupil mirror to LM1 coordinates orientation.

11:00: telescope back to zenith, CalUnit in, dichroic in, hexapod adjusted to fit sources into LGSW

DARK: 20141105_172349

Pupil PI X PI Y
Blue -0.45 -1.14
Yellow -3.38 2.15
Red -0.96 -1.75
HEXAPOD: X = -3.02, Y = -6.27, Z = 4.48, Rx = -239, Ry = 32

We center the FLAO board on the rotator using stages: X = -75, Y = -49, Z = 50

12:00: we are ready to close the loop on ARGOS, but we find a bug in the LGSW controller -- SOLVED

Alfio tells us how to solve the problem with the offload: the hexapod coordinates have to be inserted into the big PSFGUI where all the corrections must be removed and the lookup tables disabled (see HowToEnableHexapodOffload).

NOTE: at 12:30 we find a bug in the gain procedure, all the modes share the same gain value equal to the HO gain.... We will just write the HO gain in this log untill the bug has been solved...

ARGOS TN FLAO TN Notes Res.
20141105_190535 20141105_191018 ARGOS CL g =0.2, disturb =1.2" 86
20141105_191124 20141105_19452 " 93
20141105_190535-20141105_191124_modalplot.jpg

We set gain to 0.6 and we RIP the shell...

We have another problem with the RR mode that was not enabled and LB finally succeded in crashing the primary...

Gain bug fixed!

We try to move the blue pupil motor while the loop is closed without disturbance: the shell fails at 2/10 of subaperture shift.

ARGOS TN FLAO TN Notes Res.
20141105_195158 20141105_195103 OL 597
20141105_195304   " 530
20141105_195646 20141105_195529 CL, g 10 modes = 0.5, gHO = 0.4 103
20141105_195718   " 102
20141105_200024   gTT=0, Jitter OL, shell crash 145
20141105_200652   gTT=0, Jitter CL, OK 85
20141105_200740 20141105_200616 BAD: shell safe skip 119
20141105_201544 20141105_201545 g=(0.5,0.4) 99
20141105_201715   g=(0.5,0.4) 102
20141105_202052 20141105_201928 g=(0.5,0.4) 104
20141105_202656 20141105_202552 g=(0.5,0.4), FLAO TN is BAD: Alfio moved stages while acquiring 97
20141105_204541 20141105_204411 g=(0.5,0.4) 96
20141105_195158-20141105_204541_modalplot.jpg

13:20: Gilles generates a new REC from IM 20141104_195926 filtering out the TT modes from the others. Doing this we find a bug in the procedure that was filtering out TT even from TT modes themself -- SOLVED.

REC = 150modes_ryb_300Hz_TTremoved.fits

13:45: we load the new reconstructor and close again the loop on ARGOS.

ARGOS TN FLAO TN Notes Res
20141105_205114 20141105_204950 Filtered rec, g=(0.5,0.4) 91
20141105_205157   " 91
20141105_205306 20141105_205231 BAD: we opened the loop but the hexapod offload was not disabled 611
20141105_205114-20141105_205306_modalplot.jpg

There is a bug in the nice interface: when the gain is set to 0 the hexapod offload is not disabled -- SOLVED by Aflio!

We center again the LGS with the Hexapod, new PI static values:

Pupil PI X PI Y
Blue 1.00 -0.48
Yellow -1.62 0.47
Red 0.05 -0.23
We try to open again the TT putting gain=0 and to close the Jitter loop:

ARGOS TN FLAO TN Notes Res
20141105_211557 20141105_211435 Jitter CL, all other OL 600
20141105_211707   complete OL, disturb = 1.2" 600
20141105_211900 20141105_211737 Jitter CL + 10 modes on ARGOS 400
20141105_211557-20141105_211900_modalplot.jpg

14:30: we load disturb equivalent to 1.6", wind= 15m/s, 1kHz oversampling

ARGOS TN FLAO TN Notes Res
20141105_213517 20141105_213407 OL 848
20141105_213606   OL 848
20141105_213735 20141105_213610 CL, g=[0.5, 0.3] 97
20141105_213908   " 97
20141105_214031   " 99
20141105_214139   " 97
20141105_214251   " 100
20141105_213517-20141105_214251_modalplot.jpg

We load disturb equivalent to 0.8", wind= 15m/s, 1kHz oversampling

ARGOS TN FLAO TN Notes Res
20141105_214523 20141105_214359 OL 449
20141105_214607   " 448
20141105_214634   " 448
20141105_214737 20141105_214612 CL, g= [0.5, 0.3] 90
20141105_214817   " 90
20141105_214844 20141105_214727 " 90
20141105_214909   " 90
20141105_214935 20141105_214825 " 91
20141105_214523-20141105_214935_modalplot.jpg

We try now to reduce flux, until now we had 4000 ADU/subap, we go now to 2000 ADU/subap and we set a disturb equivalent to 1.2"

ARGOS TN FLAO TN Notes Res
20141105_215915 20141105_215752 2000 ADU/subap, CL g=[0.5, 0.3] 120
20141105_215942   " 119
20141105_220006   " 118
20141105_215915-20141105_220006_modalplot.jpg

We try to reduce the flux at 1000 ADU/subap and to close with gain [0.5, 0.3] but the loop diverges...

We acquire another DARK: 20141106_220741

Pupil loop seems very unstable, we try to disable it and the loop is more or less stable even with 1000 ADU/subap:

ARGOS TN FLAO TN Notes Res
20141105_221250 20141105_221159 1000 ADU/subap, CL g=[0.5, 0.3] 166
20141105_221329   " 166
20141105_221250-20141105_221329_modalplot.jpg
Topic revision: r3 - 08 Nov 2014, MarcoBonaglia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback